OpenStudy (anonymous):

another one number patterns rule thingy 17,19,22.26,31

OpenStudy (anonymous):

22. 26 thats supposts to be a camma sorry

OpenStudy (turingtest):

+2,+3,+4,+5 is the pattern so 31+6=37 37+7=44 44+8=52 52+9=61

OpenStudy (turingtest):

are the next 4 entries in the series

OpenStudy (anonymous):

? stay with me for a second

OpenStudy (turingtest):

message me if you have a question